Weight varies inversely as the distance from the center of the earth. If a woman weighs 140 pounds at sea level (4000 miles from the center of the earth), how much would she weigh in a spaceship 250 miles above the surface of the earth?
Found 2 solutions by Boreal, ikleyn:
Answer by Boreal(15235)

W=k/r
140=k/4000 units pounds*miles
k=560000
W=560000/4250 pounds*miles/miles
=131.76 pounds

Answer by ikleyn(52779)

.

The basic statement in your post is INCORRECT (as everybody knows who studied Physics, after Newton)

Weight DOES NOT vary inversely as the distance from the center of the Earth.

Weight varies inversely as the SQUARE of the distance from the center of Earth.